Bonita Companys ledger shows the following balances on December 31, 2017. 6% Preferred Stock-$10 par value, outstanding 21,500 shares Common Stock-$100 par value, outstanding 29,100 shares Retained Earnings $215,000 2,910,000 647,000 Assuming that the directors decide to declare total dividends in the amount of $398,000, determine how much each class of stock should receive under each of the conditions stated below. One years dividends are in arrears on the preferred stock. (a) The preferred stock is cumulative and fully participating. (Round the rate of participation to 4 decimal places, e.g.1.4278%. Round answers to o decimal places, e.g. $38,487.) Preferred common (b) The preferred stock is noncumulative and nonparticipating. (Round answers to o decimal places, e.g. $38,487.) Preferred Common (c) The preferred stock is noncumulative and is participating in distributions in excess of a 9% dividend rate on the common stock. (Round the rate of participation to 4 decimal places, e.g. 1.427890. Round answers to 0 decimal places, eg, $38,487.) Preferred Common
Answer #1

(a) The preferred stock is cumulative and fully participating.

Annual Dividend of preferred stock = 6% × 215000

Annual Dividend of preferred stock = $12900

Last year dividend in arrear so this year

Total dividend payable to preferred stockholder = 12900*2

Total dividend payable to preferred stockholder = 25800

Remaining Dividend = 398000 - 25800

Remaining Dividend = 372200

Total Number of share = 21500 + 29100 = 50600

Percentage of participation of common stock = 29100/50600 = 57.5099%

Percentage of participation of preferred stock = 21500/50600 = 42.4901%

Preferred Stock Dividend = $25800 + 42.4901% × $372200

Preferred Stock Dividend = $ 183948

Common Stock Dividend = 57.5099% × $372200

Common Stock Dividend = $214052

Answer

Preferred Stock Dividend = $183948

Common Stock Dividend = $214052

(b) The preferred stock is noncumulative and nonparticipating.

Annual Dividend of preferred stock = 6% × 215000

Annual Dividend of preferred stock = $12900

Since it is noncumulative and nonparticipating

Preferred Stock Dividend = $12900

Common Stock Dividend = $398000 - $12900

Common Stock Dividend = $385100

Answer

Preferred Stock Dividend = $12900

Common Stock Dividend = $385100

(c) The preferred stock is noncumulative and is participating in distributions in excess of a 9% dividend rate on the common stock.

Annual Dividend of preferred stock = 6% × 215000

Annual Dividend of preferred stock = $12900

Since it is noncumulative

Total dividend payable to preferred stockholder before participation =$12900

Total Dividend payable to common stockholder before participation = $2910000 × 9% = $261900

Remaining Dividend = $398000 - $12900 - $261900

Remaining Dividend = $123200

Total Number of share = 21500 + 29100 = 50600

Percentage of participation of common stock = 29100/50600 = 57.5099%

Percentage of participation of preferred stock = 21500/50600 = 42.4901%

Preferred Stock Dividend = $12900 + 42.4901% × $123200

Preferred Stock Dividend = $65248

Common Stock Dividend = $261900 + $123200 × 57.5099%

Common Stock Dividend = $332752

Answer

Preferred Stock Dividend = $65248

Common Stock Dividend = $332752

I have given complete and detailed explantion to your problem, in case of any query you can comment and

PLEASE DO GIVE POSITIVE RATING